# Information-Theoretic Measures on Lattices for High-Order Interactions

### v1

@@ 4 Streitberg information
It is then natural to define an information-theoretic measure based on the Möbius inversion of the full partition lattice $P(d)$ with operator function given by a divergence function.
@@
The Streitberg information $\mathrm{SI}(d)$ is defined as:
$\mathrm{SI}(d)=\sum_{\pi\in P(d)}(|\pi|-1)!(-1)^{|\pi|-1}D\left(\mathbb{P}_{\pi}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right),$ (6)
where $\mathbb{P}_{\pi}=\prod_{j=1}^{r}\mathbb{P}_{b_{j}}$ is the corresponding factorisation with respect to $\pi=b_{1}|b_{2}|\ldots|b_{r}$, and $D(\cdot\|\cdot)$ is a well-defined divergence function.
Similarly, one can define the Lancaster information:
$\mathrm{LI}(d)=\sum_{\pi_{l}\in P(d)}(-1)^{|\pi_{l}|-1}D\left(\mathbb{P}_{\pi_{l}}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)$ (7)
@@
If we choose $D(\cdot\|\cdot)$ as the KL divergence as in Eq. (5), then interaction information, Lancaster information and Streitberg information prove to be equivalent. This holds despite these measures being derived from Möbius inversions on distinct lattices.
@@
When $D(\cdot\|\cdot)$ is the KL divergence,
$\mathrm{II}(d)=\mathrm{LI}(d)=\mathrm{SI}(d)\qquad\text{for }d\geq 2$ (8)
Proof: The first equality can be proved by counting the number of double-counted marginals due to the KL formalisation. For the second equality, it suffices to show that the KL divergence associated with the non-singleton partitions (except $\hat{\mathbf{1}}$) vanish. We prove this in three ways in Section A.1 in the SI using (i) the information geometry of KL divergence, (ii) interval lattice, and (iii) the relationship between Lancaster information and Streitberg information.
The equivalence of Möbius functions for different lattices under KL divergence in Eq. (8) implies that certain interaction terms are ignored when KL is chosen. This motivates our exploring alternative divergences that exhibit different information geometries. Here, we consider Tsallis-Alpha divergence $D_{\alpha}(\mathbb{P}\|\mathbb{Q})$ [30], a classic generalisation of the KL divergence:
Definition 5 (Tsallis-Alpha divergence).
$D_{\alpha}(\mathbb{P}\|\mathbb{Q})=\frac{1}{\alpha-1}\left(\int_{x}\mathbb{P}^{\alpha}(x)\mathbb{Q}^{1-\alpha}(x)\mathrm{d}x-1\right).$ (9)
The Tsallis-Alpha divergence has a wide range of applications in machine learning, including variational inference [31], multiclass logistic regression [32] and image classification [33]. This divergence is uniquely characterised by the parameter $\alpha$: it reduces to the KL divergence when $\alpha=1$, to the Hellinger distance when $\alpha=0.5$ [34], and to the $\mathcal{X}^{2}$ distance when $\alpha=2$ [35]. The Tsallis-Alpha divergence is equivalent to the Amari-Alpha divergence up to a substitution of $\alpha$ [36].
Crucially, $D_{\alpha}(\mathbb{P}\|\mathbb{Q})$ only decomposes into lower order terms when $\alpha=1$, i.e., when it becomes KL divergence [37, 38]. As a result, $\mathrm{SI}(d)$ vanishes if the joint distribution can be factorised in any way whereas $\mathrm{LI}(d)$ vanishes if the factorisation contains at least one singleton block (Figure 3a-b). To take full advantage of the partition lattice, our focus below is $\mathrm{SI}(d)$ with Tsallis-Alpha divergence.
@@
Thus far, we have focused on the Möbius function for various lattices and operator functions, treating its counterpart, the Zeta function, primarily as an intermediate tool. However, the Zeta function leads to a recursive definition of high-order measures. Indeed, the expressions of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ can be inverted to express the divergence as a sum of $\mathrm{SI}(d)$s and $\mathrm{LI}(d)$s using the relationship in Equation (3). By rearranging the Zeta functions (see Supplementary Section A.3), we obtain
@@
$\mathrm{SI}(d)=D_{\alpha}\left(\mathbb{P}_{1\ldots d}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)-\sum_{\pi\setminus\hat{\mathbf{1}}}\mathrm{SI}(\pi),\quad\mathrm{LI}(d)=D_{\alpha}\left(\mathbb{P}_{1\ldots d}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)-\sum_{\pi_{l}\setminus\hat{\mathbf{1}}}\mathrm{LI}(\pi_{l}).$ (10)
Hence the Zeta functions of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ can be seen as attempts to recursively decompose the divergence between the joint distribution and the product of marginals. This formulation gives an intuitive interpretation when $\mathrm{SI}(d)$ is zero: the dependence in the $d$-order system can be completely explained by the lower order terms. It also follows that the sign of $\mathrm{LI}(d)$ and $\mathrm{SI}(d)$ depends on the difference between the $D_{\operatorname{KL}}\left(\mathbb{P}_{1\cdots d}\|\prod_{i=1}^{d}\mathbb{P}_{i}\right)$ and the sum of the respective low-order information.
Remark: The generalised high-order information on the partition $\pi$ can be defined using the interval lattice $[\hat{\mathbf{0}},\pi]$, e.g., $\mathrm{SI}(12|34)=D_{\alpha}(\mathbb{P}_{12}\mathbb{P}_{34}\|\mathbb{P}_{1}\mathbb{P}_{2}\mathbb{P}_{3}\mathbb{P}_{4})-D_{\alpha}(\mathbb{P}_{12}\|\mathbb{P}_{1}\mathbb{P}_{2})-D_{\alpha}(\mathbb{P}_{34}\|\mathbb{P}_{3}\mathbb{P}_{4})$. This is always zero when $\alpha=1$ due to KL decomposition hence the equivalence in Proposition 1. See Supplementary Section B for more details.
Remark: The recursiveness property underpins the definition of connected information, where it was used without explicit relation to lattice theory [14]. The terms in the connected information are elements of the powerset of $D$, making it directly comparable to $L(d)$- and $B(d)$- induced information-theoretic measures. A significant limitation of connected information is that it lacks an analytical solution; it requires the computation of the maximal entropy distribution for all elements in the powerset, thus making it less desirable than $\mathrm{SI}(d)$ defined on the complete partition lattice.
@@
Both $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ are symmetric, meaning that they are invariant under permutation of the variables. This contrasts with some information-theoretic measures such as PID that are directed, i.e., they partition the variables into a source set and a target set. In complex multivariate systems, computing a directed measure requires either looping through all the existing variables or constructing a justifiable target. Both approaches can be challenging or sometimes even unfeasible. Moreover, the decomposed information atoms are target-specific and thus insufficient for a comprehensive description of the relationships among variables.
@@
A non-zero value of $\mathrm{SI}(d)$ indicates the presence of $d$-order interaction. In particular, we say this is an emergent high-order interaction [39] if $\mathrm{SI}(d^{*})=0$ for all $12\leq d^{*}\leq d-$ 1, i.e. there are no lower order interactions. The high-order interaction can be explained solely by the dependence among $d$ variables as the only remaining term is $D(\mathbb{P}_{1\ldots d}\|\prod_{i=1}^{d}\mathbb{P}_{i})$.
Estimation of the measure
In many real-world scenarios, the distribution function of the data often remains unknown, making it impossible to directly compute information-theoretic measures. This limitation necessitates non-parametric estimators that can reliably approximate the true values of these measures. Here, we consider asymptotically unbiased estimators based on the $L_{2}$-consistent estimation of Tsallis-Alpha divergence as proposed by Poczos et al. [40] (implemented using python package ITE [41]). With this approach, we avoid direct density estimation by estimating the divergences using $k$-nearest-neighbour statistics. The resulting estimators are a linear sum of the divergences and hence consistent by Slutsky’s theorem [42].
Notice in Eq. (6) that each divergence term consists of a factorisation and its respective complete factorisation. Any factorisations that involve singletons will naturally cancel with the singletons in the complete factorisation, e.g., $D_{\alpha}(\mathbb{P}_{1}\mathbb{P}_{2}\mathbb{P}_{34}\|\mathbb{P}_{1}\mathbb{P}_{2}\mathbb{P}_{3}\mathbb{P}_{4})$ becomes $D_{\alpha}(\mathbb{P}_{34}\|\mathbb{P}_{3}\mathbb{P}_{4})$ and the cancellation makes this a 2- instead of a 4-dimensional kNN problem, leading to a more efficient computation (see Section I). As a result, summing over all partitions reduces to sum over the powerset and the partitions with no singletons due to the cancellation of the singletons. To approximate the distributions of the complete factorisation, we permute the realisations of each variables randomly to break any possible dependence among the variables, e.g., the realisations from $\mathbb{P}_{1}\mathbb{P}_{2}\mathbb{P}_{3}\mathbb{P}_{4}$ are generated by randomly permuting the realisation of from $\mathbb{P}_{1234}$.
@@ 5 Synthetic experiments
We first validate the Streitberg information using synthetic multivariate Gaussian (MVG) data (with zero mean) with ground truth interactions (analytical solution of MVG in Section D in the SI), and then investigate the accuracy of the kNN-based nonparametric estimation of the measure. Since Lancaster information and Streitberg information diverge for $d>3$ (Streitberg vanishes given any factorisation of the joint distribution while Lancaster does not due to the additional terms in the partition lattice shown in Figure 1c) we show results for $d=4$. Note that the conclusions from the below experiments also hold for $d=3,5$ and are shown in Supplementary Section E (Figures 910).
Figure 2: Robustness of Streitberg information ($d=4$). a The behaviour of $\mathrm{SI}(4)$ as a function of interaction strength is consistent across choice of $\alpha$. b Increasing the number of neighbours $k$ improves estimation accuracy. c Increasing the sample size $n$ improves estimation accuracy.
Alpha parameter
We begin by exploring the behaviour of $\mathrm{SI}(4)$ as we vary the parameter $\alpha$ using a MVG with variance $\Sigma^{1}$ where $10\leq\rho<$ 1 defines the interaction strength between variables (Figure 2a). We find varying $\alpha$ does not much change the relationship with $\rho$, highlighting the robustness of $\mathrm{SI}(d)$. With $\alpha=0.5$ the $\mathrm{SI}(4)$ is most sensitive to changes in the interaction strength and will be fixed for the below analyses.
@@
To assess the robustness of SI, we varied (i) the number of neighbours $k$ for kNN-based density estimation and (ii) the number of samples. First, we generated $n=80$ $iid$ realisations of a MVG with covariance $\Sigma^{1}$ and vary the value of $k$. Even with as few as $k=5$ neighbours, the estimation of ground truth is reasonably accurate, albeit the estimation errors when $\rho=0$ are larger for smaller $k$ (Figure 2b). Next, we fixed $k=30$ and varied the number of samples $n$. The estimations are relatively accurate even with only $n=80$ and the errors at both limits of the interaction strength are negligible. Increasing $n$ naturally improves estimation at all interaction strengths (Figure 2c).
Figure 3: Validation of Streitberg information. a $\mathrm{TC}(4)$ fails to vanish for $\mathbb{P}_{1}\mathbb{P}_{234}$, while b both $\mathrm{TC}(4)$ and $\mathrm{LI}(4)$ fail to characterise $\mathbb{P}_{12}\mathbb{P}_{34}$. $\mathrm{SI}(4)$ correctly vanishes in both cases. c The magnitude of $\mathrm{SI}(4)$ is influenced by the extent to which the joint distribution can be factorised. d Streitberg Information exhibits monotonic behaviour which is consistent across varying types of interaction.
@@
Next, we computed $\mathrm{SI}(4)$, $\mathrm{LI}(4)$ and $\mathrm{TC}(4)$ for varying analytical factorisations of the joint probability distribution: (i) $\Sigma^{2}$ where $\mathbb{P}_{1234}=\mathbb{P}_{1}\mathbb{P}_{234}$, and (ii) $\Sigma^{3}$ where $\mathbb{P}_{1234}=\mathbb{P}_{12}\mathbb{P}_{34}$. We would expect a well-behaved high-order information measure to stay at zero when a partial factorisation is present. As expected, $\mathrm{TC}(4)$ fails to capture either partial factorisation of the joint distribution, whilst $\mathrm{LI}(4)$ is only able to detect the factorisation consisting of singletons. On the other hand, $\mathrm{SI}(4)$ is able to detect both factorisations (Figure 3a-b).
Interpretation
The experiments above validate the vanishing condition of $\mathrm{SI}(d)$. Now, we illustrate the interpretation of $\mathrm{SI}(d)$ when it is non-zero. We define four different MVGs using the below covariance matrices $\Sigma^{4}$, $\Sigma^{5}$, $\Sigma^{6}$ and $\Sigma^{1}$ in Figure 3c which are equivalent when $\rho$ tends to 1 and represent block matrices with factorisable distributions when $\rho=0$. We first note that Streitberg information increases more rapidly when the covariance matrix contains more independent sub-blocks (when $\rho=0$) in Figure 3c. For covariance matrices with the same number of independent sub-blocks - such as MVG with $\Sigma^{4}$ and $\Sigma^{5}$ the latter increases more slowly. This difference can be explained by the structure of the probability distributions: pulling out one singleton from $\mathbb{P}_{12}\mathbb{P}_{34}$ creates two new singletons, whereas doing the same from $\mathbb{P}_{1}\mathbb{P}_{234}$ results in at most one new singleton. Therefore the MVG with $\Sigma^{5}$ is more amenable to further factorisation. This suggests that a larger $\mathrm{SI}(d)$ captures a lower likelihood of factorisation, pointing to stronger dependency ties within the variable set.
@@
Finally, we investigate the monotonicity of $\mathrm{SI}(d)$ using data sets constructed using an XOR gate, a COPY gate and MVG. We generate $n$ $iid$ samples of $W,X,Y,Z{\sim}\mathcal{U}(0,4)$. For the XOR gate, we set $Z_{:i}=(W_{:i}+X_{:i}+Y_{:i})\mod 4$ and $Z_{i+1:n}\sim\mathcal{U}(0,4))$. For the COPY gate, we set $W_{:i}=Z_{:i}$, $X_{:i}=Z_{:i}$, $Y_{:i}=Z_{:i}$. We then gradually increase the interaction proportion, $0\leq i/n\leq 1$. The MVG is generated with covariance matrix $\Sigma^{1}$. We find that $\mathrm{SI}(d)$ increases monotonically with interaction strength in all three examples (Figure 3d). By construction, the XOR data set does not contain pairwise or 3-way interactions, and the 4-way interaction becomes more significant as $i/n$ approaches 1. Breaking the 4-way interaction makes all the variables singletons. In contrast, the COPY and MVG data sets contain pairwise, 3-way and 4-way interactions when $i$ and $\rho$ are non-zero. In comparison to MVG, the COPY data set contains stronger interactions since the variables are eventually identical ($\rho=1$), and thus the magnitude of $\mathrm{SI}(d)$ is greater than MVG.
6 Real-world applications
Finally, we exemplify Streitberg information on two real-world datasets in finance and neuroscience.
Figure 4: $d$-order Streitberg information in stock returns within and across sectors. a The $d$-order Streitberg information was computed between stocks within sectors (coloured) and across sectors as baseline (random, grey). Each bar represents the average magnitude of $\mathrm{SI}(d)$ across $500$ samples of $d$ stocks. b We repeated the analysis for stock returns pre- and post- January 2020 (COVID onset), reporting the log ratio between the $\mathrm{SI}(d)$ within each sector and random (across sectors). Two-sample t-test within sector vs. between sector $\mathrm{SI}(d)$ were significant (*, $p<0.05$; **, $p<0.01$; ***, $p<0.001$; ****, $p<0.0001$;) except where shown as (n.s.).

### v3

@@ 4 Streitberg Information
Following our discussion above, it is then natural to define the following information-theoretic measure based on the Möbius inversion of the full partition lattice $P(d)$ with an operator function given by some divergence function acting on probability distributions.
@@
The Streitberg information $\mathrm{SI}(d)$ is then obtained from the Möbius inversion of P(d) with operator function $g(\cdot)=D(\cdot\|\prod_{i=1}^{d}p_{i})$:
$\mathrm{SI}(d)=\sum_{\pi\in P(d)}(|\pi|-1)!(-1)^{|\pi|-1}D\left(p_{\pi}\|\prod_{i=1}^{d}p_{i}\right)$ (6)
where $p_{\pi}=\prod_{j=1}^{r}p_{b_{j}}$ is the corresponding factorisation with respect to $\pi=b_{1}|b_{2}|\ldots|b_{r}$, and $D(\cdot\|\cdot)$ is a well-defined divergence between probability functions.
Similarly, one can define the Lancaster information restricted to the Lancaster (sub)lattice $L(d)$:
$\mathrm{LI}(d)=\sum_{\pi_{l}\in P(d)}(-1)^{|\pi_{l}|-1}D\left(p_{\pi_{l}}\|\prod_{i=1}^{d}p_{i}\right)$ (7)
@@
A direct consequence of our framework is that the (typical) choice of the KL divergence as the operator function leads to the equivalence of measures derived from different lattices. This is due to cancellations of interactions that lead to a collapse of the partition lattice onto particular sublattices leading to loss of information by disregarding particular interactions. In particular, it is easy to prove that, if we choose the KL divergence as the operator function $g(\cdot)=D_{\text{KL}}(\cdot\|\prod_{i=1}^{d}p_{i})$ as in Eq. (5), then interaction information, Lancaster information and Streitberg information are all equivalent, despite being derived from Möbius inversions on distinct lattices.
@@
When $D(\cdot\|\cdot)$ in Equation (6) is the KL divergence, then we have the following equivalences:
$\mathrm{II}(d)=\mathrm{LI}(d)=\mathrm{SI}(d)\qquad\text{for }d\geq 2.$ (8)
Proof.
The first equality can be proved by counting the number of double-counted marginals due to the KL formalisation. For the second equality, it suffices to show that the KL divergence associated with the non-singleton partitions (except $\hat{\mathbf{1}}$) vanish. We prove this in three ways in Appendix A.2 using (i) the information geometry of KL divergence, (ii) interval lattice, and (iii) the relationship between Lancaster information and Streitberg information. ∎
The equivalence of Möbius functions for different lattices under KL divergence shown in Equation (8) is the result of certain interaction terms cancelling out when KL is chosen as the operator function. This motivates our considering alternative divergences with different information geometry. Here, we focus on the Tsallis-Alpha divergence $D_{\alpha}(p\|q)$ [32], a classic generalisation of the KL divergence
[…]
which has found a wide range of applications in machine learning [33, 34, 35]. This divergence is uniquely characterised by the parameter $\alpha$: it reduces to the KL divergence when $\alpha=1$, to the Hellinger distance when $\alpha=0.5$ [36], and to the $\chi^{2}$ distance when $\alpha=2$ [37]. The Tsallis-Alpha divergence is equivalent to the Amari-Alpha divergence up to a substitution of $\alpha$ [38]. Crucially, $D_{\alpha}(p\|q)$ only decomposes into lower-order terms when $\alpha=1$, i.e., when it becomes the KL divergence [39, 40].
As a result, when $D(\cdot\|\cdot)$ in Equation (6) is the Tsallis-Alpha divergence, $\mathrm{SI}(d)$ vanishes if the joint distribution can be factorised in any way. In contrast, when $D(\cdot\|\cdot)$ in Equation (6) is the KL divergence, we have equivalence to $\mathrm{LI}(d)$, which vanishes if the factorisation contains at least one singleton block (Figure 3a-b). Henceforth, we focus on $\mathrm{SI}(d)$ with Tsallis-Alpha divergence, to take advantage of the factorisations in the full partition lattice.
Remark: When choosing $\alpha$ outside the range of $(0,1)$, it has been shown that the Tsallis-Alpha divergence can become unbounded [41], and similarly for the Renyi divergence, which is closely related to the Tsallis-Alpha divergence [42]. Hence we restrict our numerical experiments to the range $\alpha\in(0,1)$.
@@
Measures on lattices are generated by the Möbius function through inversion of the Zeta function. Here we show that, based on Equation (3), the expressions of $\mathrm{SI}(d)$ in Equation (6) can be expressed in terms of sums of lower-order $\mathrm{SI}(d)$, and similarly for $\mathrm{LI}(d)$ in Equation (7):
@@
$\mathrm{SI}(d)$ $=D_{\alpha}\left(p_{1\ldots d}\|\prod_{i=1}^{d}p_{i}\right)-\sum_{\pi\setminus\hat{\mathbf{1}}}\mathrm{SI}(\pi),$ (10)
$\mathrm{LI}(d)$ $=D_{\alpha}\left(p_{1\ldots d}\|\prod_{i=1}^{d}p_{i}\right)-\sum_{\pi_{l}\setminus\hat{\mathbf{1}}}\mathrm{LI}(\pi_{l}).$ (11)
Proof.
This follows from rearranging the Zeta functions (see Appendix A.3). ∎
Remark: A generalised higher-order information defined on the partition $\pi$, $\mathrm{SI}(\pi)$, can be understood using the interval lattice $[\hat{\mathbf{0}},\pi]$, e.g., $\mathrm{SI}(12|34)=D_{\alpha}(p_{12}p_{34}\|p_{1}p_{2}p_{3}p_{4})-D_{\alpha}(p_{12}\|p_{1}p_{2})-D_{\alpha}(p_{34}\|p_{3}p_{4})$. This is always zero when $\alpha=1$ due to KL decomposition, hence the equivalence in Proposition 1. See Appendix C for more details.
In this sense, the Zeta functions of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ can be seen as attempts to recursively decompose the divergence between the joint distribution and the product of marginals. This formulation gives an intuitive interpretation of the vanishing condition $\mathrm{SI}(d)=0$: the dependence in the $d$-order system can be completely explained by the lower order terms. It also follows from this rewriting that the sign of $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ depends on the difference between the $D_{\operatorname{KL}}\left(p_{1\cdots d}\|\prod_{i=1}^{d}p_{i}\right)$ and the sum of the respective lower order information terms.
Remark: The recursiveness property underpins the definition of connected information [17], where it was used without explicit links to lattice theory. The terms in the connected information are elements of the power set of $\mathcal{X}$; hence it is directly comparable to information-theoretic measures derived from the Lancaster and Boolean lattices, instead of $\mathrm{SI}(d)$, which is defined on the complete partition lattice.
@@
Both $\mathrm{SI}(d)$ and $\mathrm{LI}(d)$ are invariant under permutation of the variables. This contrasts with some information-theoretic measures such as PID [19] and PED [43] that are directed, i.e., they partition the variables into a source set and a target set. In complex multivariate systems, computing a directed measure requires either looping through all the subsets of variables or constructing a justifiable target set, both of which can be challenging or even unfeasible. Moreover, the decomposed information atoms obtained through PID are target-specific and thus insufficient for a comprehensive description of the relationships between variables.
@@
A non-zero value of $\mathrm{SI}(d)$ indicates the presence of a $d$-order interaction. A special case of this condition is when there are no lower-order interactions, i.e., if $\mathrm{SI}(d^{*})=0$ for all $12\leq d^{*}\leq d-$ 1. We denote these as emergent higher-order interactions [44], since the higher-order interaction is explained solely by the dependence among $d$ variables given that the only remaining term in Equation (10) is $D(p_{1\ldots d}\|\prod_{i=1}^{d}p_{i})$.
Consistent Estimator
In many real-world scenarios, the distribution function of the data is unknown, making it impossible to compute information-theoretic measures directly. In this scenario, non-parametric estimators that can reliably approximate the true values of our measures are needed. Here, we avoid direct density estimation and consider asymptotically unbiased estimators based on the $L_{2}$-consistent estimation of Tsallis-Alpha divergence with $k$-nearest-neighbour ($k$NN) statistics [41] (implemented using python package ITE [45]). The resulting estimators are a linear sum of the divergences and hence consistent by Slutsky’s theorem [46].
Notice in Equation (6) that each divergence term consists of a factorisation and its respective complete factorisation, but the singletons in any factorisation cancel with the singletons in the complete factorisation, e.g., $D_{\alpha}(p_{1}p_{2}p_{34}\|p_{1}p_{2}p_{3}p_{4})=D_{\alpha}(p_{34}\|p_{3}p_{4})$. This makes this latter example a 2- rather than a 4-dimensional $k$NN problem and leads to more efficient computations (see Appendix I for more computational details). As a result, summing over all partitions reduces to a sum over the power set and the partitions with no singletons. To approximate the distributions of the complete factorisation, we permute the realisations of each variable randomly to break any dependence among the variables, e.g., the realisations from $p_{1}p_{2}p_{3}p_{4}$ are generated by randomly permuting the realisation from $p_{1234}$.
Figure 2: Robustness of Streitberg information ($d=4$). All calculations in this figure are for data from a MVG with covariance matrix $\Sigma^{1}(\rho)$, where $\rho$ indicates the interaction strength. (a) The behaviour of $\mathrm{SI}(4)$ as a function of $\rho$ is consistent across $\alpha$. (b) Increasing the number of neighbours $k$ improves the estimation accuracy. (c) Increasing the sample size $n$ improves the estimation accuracy for a fixed number of neighbours ($k=5$).
@@ 5 Synthetic Experiments
To validate the Streitberg information, we use synthetic data from multivariate Gaussians (MVG) with different ground truth interactions, as defined by their correlation structure, as well as XOR and COPY gates. Since the advantages of Streitberg information become apparent for $d>3$ (see Figure 1c), we show here results for $d=4$. Similar, consistent conclusions for experiments with $d=3$ and $d=5$ are presented in Appendix D. For the analytical solution of MVG, see Appendix E.
Varying the Alpha Parameter of the Divergence
We begin by exploring the behaviour of $\mathrm{SI}(4)$ as we vary the parameter $\alpha$ of the Tsallis-Alpha Divergence $D_{\alpha}(\cdot\|\cdot)$. We use data from a MVG with variance $\Sigma^{1}$ where $10\leq\rho<$ 1 defines the interaction strength between variables (see inset of Fig. 2a). As discussed above, $\alpha\in(0,1)$ tunes the properties of the divergence with particular cases at $\alpha=0.5$ (Hellinger distance) and $\alpha=1$ (KL divergence). Our numerics in Fig. 2a show that varying $\alpha$ has little effect on the relationship with $\rho$, highlighting the robustness of $\mathrm{SI}(d)$. The $\mathrm{SI}(4)$ is most sensitive for $\alpha=0.5$, probably because $D_{\alpha}(\cdot\|\cdot)$ is symmetric in its argument only when $\alpha=0.5$. We thus fix $\alpha=0.5$ for all analyses below.
Figure 3: Validation of Streitberg information. (a) $\mathrm{TC}(4)$ fails to vanish for $p_{1}p_{234}$, while (b) both $\mathrm{TC}(4)$ and $\mathrm{LI}(4)$ fail to characterise $p_{12}p_{34}$. $\mathrm{SI}(4)$ correctly vanishes in both cases. (c) The magnitude of $\mathrm{SI}(4)$ is influenced by the extent to which the joint distribution can be factorised. (d) Streitberg Information exhibits monotonic behaviour consistently across varying types of interaction, and the magnitude indicates again the difficulty of factorising the joint distribution.
Figure 4: $d$-order Streitberg information in stock returns vary within and across sectors. (a) $\mathrm{SI}(d)$ was computed between daily returns of stocks from 2010-2024 within sectors (coloured bars) and across random sectors as baseline (grey bars) for $d=2,3,4$. Each bar represents the average magnitude of $\mathrm{SI}(d)$ across $500$ samples of $d$ stocks. Bars with deeper colours (resp. shaded) have significantly larger (resp. smaller) $\mathrm{SI}(d)$ than the inter-sector grey bars. Bars with pale colours display non-significant differences with respect to the inter-sector baseline. (b) Log ratio between the $\mathrm{SI}(d)$ within each sector and the random baseline across sectors for stock returns pre- and post-January 2020 (COVID onset). We report significance of two-sample t-test of $\mathrm{SI}(d)$ within sector vs. between sector (*, $p<0.05$; **, $p<0.01$; ***, $p<0.001$; ****, $p<0.0001$; n.s., not significant). There is an increase in significantly different values of $\mathrm{SI}(4)$ post-COVID across sectors.
@@
To assess the robustness of the estimation of Streitberg information, we varied (i) the number of neighbours $k$ for $k$NN-based density estimation and (ii) the number of samples $n$. First, we generated $n=80$ $iid$ realisations of a MVG with covariance $\Sigma^{1}$ and varied the value of $k$. Even with as few as $k=5$ neighbours, the estimation of the ground truth is reasonably accurate, albeit with larger estimation errors for $\rho=0$ (Figure 2b). Next, we fixed $k=30$ and varied the number of samples $n$. The estimations are relatively accurate even with only $n=80$ with improved estimation for all $\rho$ as $n$ is increased (Figure 2c).
@@
Next, we computed $\mathrm{SI}(4)$, $\mathrm{LI}(4)$ and $\mathrm{TC}(4)$ for varying factorisations of $p_{1234}$: (i) $\Sigma^{2}$ where $p_{1234}=p_{1}p_{234}$, and (ii) $\Sigma^{3}$ where $p_{1234}=p_{12}p_{34}$. A well-behaved higher-order information-theoretic measure should stay at zero when a lower-order factorisation of the joint distribution is present. As expected, $\mathrm{TC}(4)$ fails to capture either factorisation, whilst $\mathrm{LI}(4)$ is only able to detect the factorisation consisting of singletons. Importantly, $\mathrm{SI}(4)$ is able to detect both factorisations (Figure 3a-b).
Interpretation of $\mathrm{SI}(d)$
Now, we illustrate the interpretation of $\mathrm{SI}(d)$ when it is non-zero. We define four different MVGs using the covariance matrices $\Sigma^{4}$, $\Sigma^{5}$, $\Sigma^{6}$ and $\Sigma^{1}$ in Figure 3c which are equivalent when $\rho$ tends to 1 and represent block matrices with factorisable distributions when $\rho=0$. Figure 3c shows that $\mathrm{SI}(4)$ increases more rapidly as a function of the interaction strength $\rho$ when the covariance matrix contains more independent sub-blocks (in the factorisable limit of $\rho=0$). For covariance matrices $\Sigma^{4}$ and $\Sigma^{5}$, which have the same number of factorisable sub-blocks in the limit, $\mathrm{SI}(4)$ increases more slowly for $\Sigma^{5}$. This difference can be explained by the structure of the probability distributions: pulling out one singleton from $p_{12}p_{34}$ creates two new singletons, whereas doing the same from $p_{1}p_{234}$ results in at most one new singleton. Therefore the MVG with $\Sigma^{5}$ is more amenable to further factorisation. This suggests that a larger $\mathrm{SI}(d)$ captures a lower likelihood of factorisation, pointing to stronger dependency ties within the set of variables.
@@
Finally, we investigate the monotonicity of $\mathrm{SI}(d)$ using data sets from an XOR gate, a COPY gate and MVG. The MVG is generated with covariance matrix $\Sigma^{1}$. For the XOR and COPY gates, we proceed as follows. We generate $n$ $iid$ samples of $W,X,Y,Z{\sim}\mathcal{U}(0,4)$. For the XOR gate, we set $Z_{:i}=(W_{:i}+X_{:i}+Y_{:i})\mod 4$ and $Z_{i+1:n}\sim\mathcal{U}(0,4))$. For the COPY gate, we set $W_{:i}=Z_{:i}$, $X_{:i}=Z_{:i}$, $Y_{:i}=Z_{:i}$. We then gradually increase the interaction proportion, $0\leq i/n\leq 1$. We find that $\mathrm{SI}(d)$ increases monotonically with the interaction strength in all three examples (Figure 3d). By construction, the XOR data set does not contain pairwise or 3-way interactions, and breaking the 4-way interaction makes all the variables into singletons. In contrast, the COPY and MVG data sets contain pairwise, 3- and 4-way interactions when $i$ and $\rho$ are non-zero. In comparison to MVG, the COPY data set contains stronger interactions since the variables are eventually identical ($\rho=1$), and thus the magnitude of $\mathrm{SI}(d)$ is greater than MVG.
6 Applications
Finally, we exemplify $\mathrm{SI}(d)$ on applications in finance, neuroscience and machine learning.